# Latent Assistance Networks: Rediscovering Hyperbolic Tangents in RL

### v1

@@ 4 Experiments
We conduct our main experiments and discussion on the Deep Q-Learning algorithm (DQN) by Mnih et al. (2015) in Atari, and complement them by additionally evaluating the LaN on the Proximal Policy Optimization algorithm (PPO) (Schulman et al., 2017). In all experiments, we build on the strong baseline implementations from cleanrl  (Huang et al., 2022). For both DQN and PPO, we define the representation $z_{t}$ as the last hidden layer of the network, which serves as the compressed representation of the original pixel observation $s_{t}$ (see Fig. 3). Using a LaN, we show that it is possible to decrease dying hyperbolic tangents while respectively increasing the effective rank of the representation $z_{t}$, as well as downstream performance.
@@
Figure 5: Normalized performance with the standard deviation over the means in the Atari domain, after training DQN for 10M iterations (40M Frames). In (a), a product of hyperbolic tangents is the only continuously differentiable activation that performs equal or better than the ReLU activation, without any hyperparameter changes. Furthermore, using a product of hyperbolic tangent representations seems to greatly outperform the single hyperbolic tangent. In (b), several ablations including layer normalization (Ba et al., 2016) on hyperbolic tangent activations are shown. Notably, the decrease in performance when using a single hyperbolic tangent with a twice as large representation ($z_{t}\in\mathbb{R}^{1024}$) shows that more parameters are not always desirable.
We perform all experiments on 8 common, non-exploration driven Atari games. We normalize using the ReLU baseline normalized scores and show in Appendix D.1 that human-normalized scores provide an even better score for Tanh (LaN), due to the strong LaN performance in one of the games where humans perform much worse.
Reinforcement Learning
$\mathcal{L}_{Q}=\big{|}Y_{t}-Q(z_{t},a)\big{|}^{2}.$ (4)
4.1 Atari - DQN
(a) Dead Neurons
@@ 4.1 Atari - DQN
Figure 6: The average fraction of dead neurons (a) and the average effective rank (Kumar et al., 2021) (b) of the representation $z_{t}$ when training DQN in the Atari domain for 10M iterations (40M frames). Similar to the well-known dying ReLU problem, hyperbolic tangent and sigmoid activations also exhibit strong dying neuron behavior when training RL. Furthermore, a product of hyperbolic tangents reduces dead neurons in $z_{t}$ and subsequently increases the effective rank of the representation.
Figure 7: Additional ablations run on DQN in the Atari domain. Tanh (+LaN) represents an addition rather than an elementwise product with a LaN, and Tanh (2xLaN) uses two Latent assistance Networks.
The influence of a LaN on downstream performance can be seen in Fig. 5. A significant improvement over the standard hyperbolic tangent baseline is seen, as well as an important improvement over the strong ReLU baseline. Examining further ablations in Fig. 7 shows that using a sophisticated Linear-Unit function such as the SELU activation (Klambauer et al., 2017) or using an addition rather than a product of hyperbolic tangents seems detrimental to performance. Furthermore, taking a product of 3 hyperbolic tangents (2$\times$×LaN) also appears to enhance performance, though the improvement is not significantly different from using a single LaN. This could be the result of increased contracting behavior in the early stage of training due to subsequent multiplication of activations whose absolute values are $<1$. Further experiments combining the ReLU activation with a LaN can be found in Appendix C.3.
Mitigating dying Hyperbolic Tangents
As mentioned in Section 3, the number of dead neurons is defined as the amount of neurons that display the same output for any given observation $s_{t}$. The amount of dead neurons over time when training on Atari can be seen in  6(a). For the ReLU activation, around 60% of the neurons in the representation $z_{t}$ are dead, while for the sigmoid and hyperbolic tangent activation this number is around 40%. When using a hyperbolic tangent with a LaN, a reduction in dead neurons as compared to using a single hyperbolic tangent or sigmoid can be observed. We credit this to the inherent ability of a product of hyperbolic tangents to minimize long-term activation saturation, as explained in Section 3. This reduction of long-term dead hyperbolic tangents when using the LaN is remarkable as it is also coinciding with a stronger divergence of values in the initial stages of training (see Fig. 4).
Increasing Effective Rank
We calculate the effective rank (Kumar et al., 2021) of the representation $z_{t}$ during training, which can be seen in Fig. 6(b). As observed by Gulcehre et al. (2022), a representation activated by a hyperbolic tangent or, residing in the same category, a sigmoid, already has a relatively high effective rank compared to a representation activated by a ReLU. Furthermore, as can additionally be seen in supervised learning (see Fig. 2), a LaN notably improves the effective rank of the hyperbolic tangent, which can be seen as a measure of network ‘expressivity’. We summarize our results qualitatively in Table 2.
Table 2: High-level summary of the strengths of a LaN.
Activation function Performance RL Few dead neurons Effective rank
Tanh (LaN) ++ ++ ++
@@
4.2 Atari - PPO
Additional experiments are run using the PPO algorithm (Schulman et al., 2017). For PPO, the internal architectural difference with DQN is that the compressed representation $z_{t}$ precedes both a critic and an actor network, and thus receives policy and value gradients (see Appendix A.2). The results can be found in Fig. 8. Similarly to DQN, the Tanh (LaN) exhibits the highest effective rank and the best performance. The Tanh (LaN) and the ReLU also seem to have consistent strong performance across both algorithms, whereas the pure hyperbolic tangent and sigmoid activations are relatively unreliable.
(a) Normalized Score
(b) Effective Rank
@@ 5 Related Work
Liu et al. (2019) investigated the need for sparse representations in the continuous control domain. Gulcehre et al. (2022) analyzed network expressiveness in RL by measuring the effective rank (Kumar et al., 2021) of the compressed representation, and found that hyperbolic tangent representations generally maintain high rank and do not suffer strongly from rank decay as training continues. Other work by Lyle et al. (2022) investigated capacity loss in RL and similarly found that, as training progresses, the inherent network capacity of RL algorithms decays. Further research by Nikishin et al. (2022) used network resets to counteract the primacy bias and Sokar et al. (2023) evaluated and mitigated the dying ReLU phenomenon in DQN, both operating in the sample efficiency setting. Nikishin et al. (2023) further studied plasticity injection for long-term training. In another related direction, recent work has investigated network sparsity in RL, showing that a large part of network capacity might be unnecessary when training reinforcement learning (Graesser et al., 2022; Sokar et al., 2022; Tan et al., 2023; Arnob et al., 2021). This provides further insights into why a ReLU can achieve strong performance despite resulting in a significant number of dead neurons.
Network Architecture
The origin of network optimization problems with hyperbolic tangents and sigmoids were empirically investigated by Glorot & Bengio (2010), where, according to the authors, a lot of mystery still surrounds the subject. Work by Srivastava et al. (2015) in supervised learning first looked at the idea of using products of hidden layers together with a ’gate’ that determined the amount of information flow (Hochreiter & Schmidhuber, 1997). Using these ideas, the Resnet was invented (He et al., 2016). More specific to RL, Henderson et al. (2018) showed differences in RL performances over different network architectures and nonlinear activations. Work by Abbas et al. (2023) successfully applied ReLU concatenation (Shang et al., 2016) to improve continual learning while keeping a similar performance when training from scratch. Finally, recent work by Grooten et al. (2024) investigated raw pixel masking for distractions in RL using a parallel CNN input layer.
@@ 6 Conclusions and Discussion
We have proposed a novel architecture called the Latent assistance Network (LaN) that augments a networks’ final hidden layer by taking the element-wise product with a parallel, independent hidden layer. We discussed and empirically showed that this decreases the occurrence of dead neurons in the representation $z_{t}$, increases network expressiveness, and notably improves performance compared to a single continuously differentiable activation such as the hyperbolic tangent.
Our results also show an important increase in performance over the dominant ReLU activation. This paper therefore remains, to the best of our knowledge, the first to show the possibility of elevating a continuously differentiable activation to and beyond the level of the ReLU. Additionally, it should be recognized that the employed DQN baseline was initially optimized specifically for the ReLU activation function.
As investigated by Kumar et al. (2021), a compressed representation activated by a hyperbolic tangent has a much higher effective rank than the ReLU and could therefore yield stronger network expressiveness. We therefore remain hopeful that the hyperbolic tangent can potentially replace the ReLU in a variety of architectures. Making strong claims in this field however remains challenging, as the success of pruning strategies show that network expressiveness might already be in abundance in some settings (Frankle & Carbin, 2019; Obando-Ceron et al., 2024).
Future work could focus on further identifying the intricacies of this architecture in an attempt to push the potential of continuously differentiable activations in reinforcement learning or even in supervised learning. Also, we believe that an implementation of this architecture in a continual learning setting with similar evaluations as the work in Abbas et al. (2023) could be promising.
@@ 7 Limitations
Our work has focused on the encoders’ final hidden layer in reinforcement learning. As literature suggests that convolutional layers have an even stronger preference for ReLU activations (Jarrett et al., 2009), we have not conducted experiments testing a LaN on the convolutional activations of our encoders.
Due to the heavy computational requirements of the Atari domain, limitations of our work include the use of a moderate amount of 5 seeds per environment, as well as the lack of a hyperparameter search for using the ‘Tanh (LaN)’ setting in our experiment section. A hyperparameter search could further strengthen our results, as the strong baseline is specifically tuned for the ReLU activation. Furthermore, as we perform experiments on DQN and PPO, integration of our Latent assistance Network into more sophisticated algorithms such as Rainbow (Hessel et al., 2018) would be interesting.
@@ A.2 PPO Architecture
Figure 9: A visualisation of the Latent assistance Network (LaN) architecture combined with the PPO architecture applied on a snapshot of the ’SpaceInvaders’ Atari environment. A parallel linear layer providing an independent representation is integrated, where the element-wise product of the two parallel representations represents the final representation $z_{t}$.
@@ Appendix B Kernel Density Estimations
As discussed in Section 3, we hypothesize that the differences between a hyperbolic tangent with and without a LaN are due to the increased ability of the product of hyperbolic tangents being able to negate dying neurons. We further see this phenomenom when plotting a random selection of neurons from both the mask and the base representation in Fig. 10(a).
(a) Tanh (LaN) - $5\cdot 10^{6}$ iterations
(b) Tanh (LaN) - $5\cdot 10^{6}$ iterations
(c) Tanh - $5\cdot 10^{6}$ iterations
Figure 10: Kernel Density Estimations (KDE) over a subset of 16 neurons in the representations $z_{t}^{enc}$ and $z^{*}_{t}$ in (a), the resulting element-wise product $z_{t}$ in (b) and the representation $z_{t}$ when training without a LaN (c). These representations are obtained after training DQN in the ’Breakout’ environment. A closer look at neurons 3, 8 and 9 show that when one of the representations saturates, the other is able to compensate, leading to a non-dead neuron in their product $z_{t}$.
@@
(d) Sigmoid (LaN) - $10^{7}$ iterations
Figure 11: Kernel Density Estimations of the final representation $z_{t}$ after training DQN for $10^{7}$ iterations in the Breakout environment. A hyperbolic tangent using LaN allows the representation to avoid strong saturation, keeping sufficient kernel density in the central sections of the hyperbolic tangent. As a sigmoid can saturate into zero, using a LaN remains less effective into preventing saturation.

### v2

@@ 4 Experiments
The main experiments and discussion are conducted with the value-based DQN algorithm (Mnih et al., 2015) in Atari, and complemented by an evaluation on the policy-based proximal policy optimization algorithm (PPO) (Schulman et al., 2017). In all experiments, we build on the implementations from cleanrl  (Huang et al., 2022). For more details on the PPO and DQN implementation, we refer the reader to Appendix A.1. For both DQN and PPO, we define the representation $z_{t}\in\mathbb{R}^{512}$ as the last hidden layer of the network, which serves as the compressed representation of the original pixel observation $s_{t}$ (see Fig. 4).
We perform all experiments on 8 common, non-exploration driven Atari games. We aim to show the effect of a Hadamard representation on the representation’s fraction of dying neurons, its effective rank and its downstream performance. For more information on the dying neuron and effective rank calculations, we refer the reader to Appendix. B.1 and Appendix. B.2, respectively. The performance is aggregated over 8 environments and normalized using the ReLU baseline scores.
(a) Dead Neurons
(b) Effective Rank
Figure 6: The average fraction of dead neurons (a) and the average effective rank (Kumar et al., 2021) (b) of the representation $z_{t}$ when training DQN in the Atari domain for 10M iterations (40M frames). Similar to the well-known dying ReLU problem, hyperbolic tangent and sigmoid activations also exhibit strong dying neuron behavior during training. A Hadamard product of hyperbolic tangents reduces dead neurons in $z_{t}$ and subsequently increases the effective rank of the representation.
Mitigating dying Hyperbolic Tangents
As mentioned in Section 3 and defined in Eq. 2, the number of dead neurons is equal to the amount of neurons that display the same saturated output for any given observation $s_{t}$. The amount of dead neurons over time when training on Atari can be seen in Fig.  6(a). For the ReLU activation, around 60% of the neurons in the representation $z_{t}$ die during training, while for the sigmoid and hyperbolic tangent activation this number is around 40%. When using a hyperbolic tangent Hadamard representation, a reduction in dead neurons as compared to using a single hyperbolic tangent can be observed. We credit this to the inherent ability of a Hadamard product of hyperbolic tangents to minimize long-term activation saturation, as explained in Section 3. Quantitative results of dead neurons can be found in Table 2, which is correlated with earlier predictions from Table 1.
Table 2: Dying neuron fractions with and without a Hadamard representation (HR).
Activation Function Dead Neuron Fraction Dead Neuron Fraction (HR) Difference
Hyperbolic Tangent 0.39 0.30 -23%
Sigmoid 0.44 0.45 +2%
ReLU 0.62 0.73 +18%
For more dying neuron graphs related to the ReLU and the Sigmoid activations, we refer the reader to Appendix C.3.
Increasing Effective Rank
We additionally investigate the effective rank (Kumar et al., 2021) of the representation $z_{t}$ during training, which can be seen in Fig. 6(b). As observed by Gulcehre et al. (2022), a representation activated by a hyperbolic tangent or a sigmoid, already has a relatively high effective rank compared to a representation activated by a ReLU. Furthermore, similar to the results in our supervised learning experiments (see Fig. 2), using a Hadamard representation with hyperbolic tangents significantly improves the effective rank, which is strongly correlated to a network’s ‘expressivity’. Employing a Hadamard representation with ReLU activations significantly decreased the effective rank of the representation, as was expected. For more information on the effect of a Hadamard representation on a ReLU, we refer the reader to Appendix C.4.
Analyzing Performance in Atari
The influence of a Hadamard representation on downstream performance can be seen in Fig. 7. Correlating with the reduction in dying neurons and an increase in effective rank, a significant improvement of nearly $60\%$ over the standard hyperbolic tangent baseline is obtained, as well as an improvement over the default ReLU baseline. Note that no hyperparameter or architectural tuning has been employed in our experiments.
@@
Figure 7: Normalized performance with the standard deviation over the means in the Atari domain, after training DQN for 10M iterations (40M Frames). In (a), the Tanh (HR) significantly outperforms a the baseline Tanh. In (b), several ablations including layer normalization (Ba et al., 2016) on hyperbolic tangent activations are shown. Notably, the decrease in performance when using a conventional representation with twice as much neurons ($z_{t}\in\mathbb{R}^{1024}$) shows that more parameters are not always desirable. Additional experiments combining the ReLU activation with a Hadamard representation can be found in Appendix C.4.
Figure 8: Additional ablations run on DQN in the Atari domain. Tanh (+) represents an addition rather than a Hadamard product, and Tanh (2HR) uses a triple Hadamard product.
Examining further ablations in Fig. 8 shows that using a sophisticated Linear-Unit function such as the SELU activation (Klambauer et al., 2017) or using an addition rather than a product of hyperbolic tangents seems detrimental to performance. Furthermore, taking a product of 3 hyperbolic tangents (2HR) also appears to enhance performance, though there seems to be a negative effect in the early stages of training as compared to using a single Hadamard product. We hypothesize that this is the result of increased contracting behavior in the early stage of training due to increasing multiplication of hyperbolic tangent activations whose absolute values are $<1$. For human-normalized scores, we refer the reader to Appendix D.1.
Evaluating Hadamard representations on PPO
To evaluate results on a policy-based algorithm, additional experiments are run using the PPO algorithm (Schulman et al., 2017). For PPO, the internal architectural difference with DQN is that the compressed representation $z_{t}$ precedes both a critic and an actor network, and thus receives policy and value gradients (see Appendix A.2). The results can be found in Fig. 9. Similarly to the results on DQN, the Tanh (HR) exhibits the highest effective rank and the best performance. The Tanh (HR) and the ReLU also seem to have consistent strong performance across both algorithms, whereas the pure hyperbolic tangent and sigmoid activations can be relatively unreliable across algorithms.
(a) Normalized Score
@@
Interestingly, it seems that the ReLU activations’ performance is less correlated to its low effective rank and high amount of dying neurons than the continuously differentiable activations. As also indicated by Teney et al. (2024), the exact reasons for the broad success of the ReLU activations are not yet fully understood. A qualitative overview of our results can be found in Table 3.
Table 3: High-level summary of the strengths of a Hadamard representation.
Activation function Performance Few dead neurons Effective rank
Tanh (HR) ++ ++ ++
@@
Increasing Representation Dimensionality
A Hadamard representation increases the amount of trainable parameters in the encoder without increasing the dimension of the representation $z_{t}\in\mathbb{R}^{512}$. This is an important feature as simply increasing the dimension of the representation to $\mathbb{R}^{1024}$ is not often effective (see Fig. 7(b)). Even though the fraction of dying neurons decreases and the effective rank increases when using $z_{t}\in\mathbb{R}^{1024}$, it seems that the ’relative’ effective rank, defined as the ratio $\frac{rank}{dim(z_{t})}$, actually goes down (See Appendix. C.2). We therefore hypothesize that reducing dying neurons and increasing a representation’s effective rank is only superficially solved by expanding representation dimensionality.
@@ 5 Related Work
Liu et al. (2019) investigated the need for sparse representations in the continuous control domain. Gulcehre et al. (2022) analyzed network expressiveness in RL by measuring the effective rank (Kumar et al., 2021) of the compressed representation, and found that hyperbolic tangent representations generally maintain high rank while not suffering strongly from rank decay as training continues. Related work used normalization techniques and action penalization to counteract high variance in pixel-based robotic control (Bjorck et al., 2022). Other work by Lyle et al. (2022) investigated capacity loss in RL and similarly found that, as training progresses, the inherent network capacity of RL algorithms decays. Further research by Nikishin et al. (2022) used network resets to counteract the primacy bias and Sokar et al. (2023) evaluated and mitigated the dying ReLU phenomenon in DQN, both operating in the sample efficiency setting. Nikishin et al. (2023) further studied plasticity injection for long-term training. In another related direction, recent work has investigated network sparsity in RL, showing that a large part of network capacity might be unnecessary when training reinforcement learning (Arnob et al., 2021; Graesser et al., 2022; Sokar et al., 2022; Tan et al., 2023; Obando-Ceron et al., 2024). This provides further insights into why a ReLU can achieve strong performance despite resulting in a significant number of dead neurons.
Network Architecture in RL
The origin of network optimization problems with hyperbolic tangents and sigmoids were empirically investigated by Glorot & Bengio (2010), where, according to the authors, a lot of mystery still surrounds the subject. Work by Srivastava et al. (2015) in supervised learning first looked at the idea of using products of hidden layers together with a ’gate’ that determined the amount of information flow (Hochreiter & Schmidhuber, 1997). Using these ideas, the Resnet was invented (He et al., 2016) and also showed strong performance in combination with RL (Espeholt et al., 2018). Further work by Henderson et al. (2018) showed differences in RL performances over different network architectures and nonlinear activations. Work by Abbas et al. (2023) successfully applied ReLU concatenation (Shang et al., 2016) to improve continual learning while keeping a similar performance when training from scratch. Finally, recent work by Grooten et al. (2024) investigated raw pixel masking for distractions in RL using a parallel CNN input layer.
@@ 6 Conclusions and Discussion
This paper analyzed issues with continuously differentiable activations in RL and demonstrated that these activation functions also suffer from the dying neuron problem. Based on this analysis, we propose a novel representation architecture, the Hadamard Representation (HR), which enhances an encoder’s final hidden layer by taking the Hadamard product with a parallel, independently parameterized activation layer. We further discussed and empirically showed that applying the Hadamard representation with hyperbolic tangents reduces the occurrence of dead neurons in the representation and increases layer expressiveness. This approach significantly improved performance in Atari games compared to both standard representation parameterizations and merely increasing the representation dimension.
As investigated by Kumar et al. (2021), a compressed representation activated by a hyperbolic tangent attains a much higher effective rank than when activated by a ReLU, and will therefore yield stronger network expressiveness. In light of this, the hyperbolic tangent can potentially replace the ReLU in a variety of architectures where expressiveness is required without using deeper networks.
Future work could focus on further identifying the intricacies of hyperbolic tangents in an attempt to push the potential of continuously differentiable activations in reinforcement learning or even in a supervised learning setting (see Fig. 2). Also, we believe that an implementation of Hadamard-style architectures in a continual learning setting as in Abbas et al. (2023) could be promising.
@@ 7 Limitations
Our research has focused on the encoders’ final hidden layer in reinforcement learning. Both literature and preliminary experiments suggest that convolutional layers have a very strong preference for rectified activations (Jarrett et al., 2009), possibly due to a low-frequency bias (Ramasinghe & Lucey, 2022; Teney et al., 2024). We have therefore not conducted experiments integrating Hadamard representations in the convolutional section of our encoder architectures. In addition, a hyperparameter search could further strengthen our results, since the baseline is specifically tuned for the ReLU activation. Furthermore, as we perform experiments on DQN and PPO, integration of a Hadamard representation into more sophisticated algorithms and architectures such as Rainbow (Hessel et al., 2018) and Impala (Espeholt et al., 2018) would be interesting.
Reproducibility Statement
Constructing a Hadamard representation is a relatively straightforward process, of which the Pytorch pseudocode can be found in Appendix A.1. The DQN and PPO algorithms used are adopted from the compact cleanrl baselines (Huang et al., 2022). Finally, the details on the KDE calculation and the effective rank calculation can be found in Appendix B.1 and Appendix B.2.
@@ A.1 Hyperparameters
Hadamard Implementation
Constructing a Hadamard representation is a straightforward process that only requires an additional, parallel linear layer. Starting from the flattened convolutional features in DQN, the Pytorch pseudocode is defined as follows:
flattened_features = conv_features.flatten(1)
representation1 = nn.Tanh(linear1(flattened_features))
representation2 = nn.Tanh(linear2(flattened_features))
hadamard = representation1 * representation2
Q-values = linear_q(hadamard)
Reinforcement Learning
$\mathcal{L}_{Q}=\big{|}Y_{t}-Q(z_{t},a)\big{|}^{2}.$ (4)
@@ A.2 PPO Architecture
Figure 10: A visualisation of the Hadamard representation (HR) architecture combined with the PPO architecture applied on a snapshot of the ’SpaceInvaders’ Atari environment. A parallel linear layer providing an independent representation is integrated, where the Hadamard product of the two parallel representations represents the final representation $z_{t}$.
@@ Appendix B Kernel Density Estimations
As discussed in Section 3, we hypothesize that the differences between a hyperbolic tangent with and without an HR are due to the increased ability of the product of hyperbolic tangents being able to negate dying neurons. We further see this phenomenom when plotting a random selection of neurons from both the mask and the base representation in Fig. 11(a).
(a) Tanh (HR) - $5\cdot 10^{6}$ iterations
(b) Final Tanh (HR) - $5\cdot 10^{6}$ iterations
(c) Tanh (no HR) - $5\cdot 10^{6}$ iterations
Figure 11: Kernel Density Estimations (KDE) over a subset of 16 neurons in the representations $z_{t}^{enc}$ and $z^{*}_{t}$ in (a), the resulting Hadamard product $z_{t}$ in (b) and the representation $z_{t}$ when training without an HR (c). These representations are obtained after training DQN in the ’Breakout’ environment. Red outlines represent dead (collapsed) neurons. In (a), a closer look at neurons 3, 8 and 9 shows that when one of the representations saturates, the other is able to compensate, leading to a non-dead neuron in their product $z_{t}$ in (b).
@@
(d) Sigmoid (HR) - $10^{7}$ iterations
Figure 12: Kernel Density Estimations of the final representation $z_{t}$ after training DQN for $10^{7}$ iterations in the Breakout environment. A hyperbolic tangent Hadamard representation allows the representation to avoid strong saturation, keeping sufficient kernel density in the central sections of the hyperbolic tangent. As a sigmoid can saturate into zero, using a Hadamard representation remains less effective for preventing saturation, as any zero will lead to a Hadamard product of zero.
@@ B.1 KDE calculation
B.2 Effective Rank calculation
In line with Kumar et al. (2021), the effective rank of a feature matrix for a threshold $\delta$ ($\delta=0.01$), denoted as $srank_{\delta}(\Phi)$, is given by $srank_{\delta}(\Phi)=\min\left\{k:\frac{\sum_{i=1}^{k}\sigma_{i}(\Phi)}{\sum_{i=1}^{d}\sigma_{i}(\Phi)}\geq 1-\delta\right\}$, where $\{\sigma_{i}(\Phi)\}$ are the singular values of $\Phi$ in decreasing order, i.e., $\sigma_{1}\geq\cdots\geq\sigma_{d}\geq 0$. Intuitively, the effective rank of a feature matrix represents the number of “effective” unique components that form the basis for linearly approximating the resulting Q-values. The calculation in Python is done as follows:
def compute_rank_from_features(feature_matrix, rank_delta=0.01):
    sing_values = np.linalg.svd(feature_matrix, compute_uv=False)
    cumsum = np.cumsum(sing_values)
    nuclear_norm = np.sum(sing_values)
    approximate_rank_threshold = 1.0 - rank_delta
    threshold_crossed = (cumsum >= approximate_rank_threshold * nuclear_norm)
    effective_rank = sing_values.shape[0] - np.sum(threshold_crossed) + 1
    return effective_rank